OpenStudy (anonymous):

find the value of a that makes the system inconsistent -3x+5y=2 ax-10y=0

OpenStudy (jiteshmeghwal9):

find the value of x from one of the equation as\[-3x=2-5y \implies x=\frac{2-5y}{-3}\]

OpenStudy (jiteshmeghwal9):

now plug the value of x in the second equation as\[a{{(2-5y)} \over {(-3)}}-10y=0\]

OpenStudy (jiteshmeghwal9):

now solve it for a:)

OpenStudy (anonymous):

what do i do from a (2-5y)=-30y

OpenStudy (jiteshmeghwal9):